# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ 3.1. Definition
Federated Learning (Society, 2021; McMahan et al., 2017) is a collaborative machine learning modeling paradigm that enables sharing and aggregation of knowledge from multiple sources while maintaining the confidentiality of source data. Generally, FL systems consist of two main entities in terms of task organization: the server and the participants. The FL server can launche a federated training task and invites participants with sufficient training data and hardware resources to contribute their local modeling results for multi-source knowledge aggregation. In practice, FL systems can be divided into two categories based on application scenarios (Kairouz et al., 2021):
  • Cross-device FL. In this setting, the participants are numberous end devices with relatively small dataset size, such as mobiles, IoT sensors and wearable devices, the server is hosted in the cloud. Since there is low context correlation between the data of distributed end devices and less overlapping sample ids, this setting typically falls within the scope of horizontal FL. The cross-device FL applications include: Gboard input suggestion (Hard et al., 2018; Ramaswamy et al., 2019; Yang et al., 2018), e-commerce recommendation (Niu et al., 2020) and advertising (Wang et al., 2023a).
  • Cross-silo FL. In this setting, the participants are orginizations or institutions with large amounts of well-maintained structured data, and the server is hosted by a trusted FL service providers such as FATE (Liu et al., 2021b) and NVFLARE (Roth et al., 2022). As participants can be different departments within an organization, the data silo owned by these departments can have a large overlap in sample space and less overlap in feature space, which falls within vertical FL. The applications of cross-silo FL include federated data analysis for radiomics (Li et al., 2019, 2020a; Scherer et al., 2020), epidemiology (Dayan et al., 2021) and EHR (Brisimi et al., 2018; Huang et al., 2019).
@@
Furthermore, FL entities can also serve multiple functional roles to support advanced features such as privacy enhancement (Geyer et al., 2017; Bonawitz et al., 2017; Niu et al., 2020), participant scheduling (Li et al., 2022d; AbdulRahman et al., 2020a), model verification (Tekgul et al., 2021; Shao et al., 2022) and incentive mechanisms (Yu et al., 2020). Recall that there are four roles defined in the FL standard (Society, 2021):
@@
  • Model User. The FL model users can request for FL modeling services and preset the targeted task, and then establish cooperation with participants who provide training data. This role can leverage the benefits of collaborative training to improve the preformance of its objective models.
@@
  • Coordinator. The FL coordinators are responsible for providing FL services to all FL entities. This role involves setting up communication channels with entities, initializing the execution environment of participants (Hanzlik et al., 2021), scheduling the training and aggregation workflows for improve system efficiency, such as by alleviating the straggler effect (Li et al., 2021a; Chai et al., 2020), optimizing data heterogeneity (Duan et al., 2019; AbdulRahman et al., 2020a) and compressing model transfer (Konečnỳ et al., 2016; Sattler et al., 2019). Additionaly, the FL coordinator provides privacy control mechanisms (Bonawitz et al., 2017; El Ouadrhiri and Abdelhadi, 2022; Hesamifard et al., 2018) for model users and authorization verification for participants to maintain the security of FL systems. Furthermore, the coordinator can hold a validation dataset for evaluate the models contributed by participants or detect potential disturbances from Byzantine attacks (Sattler et al., 2020).
@@
  • Data Owner. The FL data owners are knowledge contributors of FL systems, they collect and desentize raw data to maintain a local dataset for federated training. Although they have full authority of data processing and modeling, they cannot share the raw data due to privacy concerns. To address these concerns, de-identification (Act, 1996) and differential privacy (Dwork, 2006) techniques can be applied to meet privacy budgets as required by privacy policies.
@@
  • Auditor. The FL auditors are responsible for formulating privacy control policies and establishing supervisory mechanisms that ensure the training process is compliant with data protection regulations (e.g. HIPAA (Act, 1996), GDPR (Voigt and Von dem Bussche, 2017)) and preventing potential privacy breaches for both model users and data owners. Especially in FL, the latent knowledge in models can potentially reveal the sensitive information of training data (Wang et al., 2019; Zhu et al., 2019; Jin et al., 2021), making it crucial for auditors to scrutinize the model transmission (Wei et al., 2021; Li et al., 2022f) and verify the ownership of models (Tekgul et al., 2021; Shao et al., 2022).
@@
Figure 2. An overview of traditional FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
[…]
@@
Fig. 2 illustrates the typical architecture of FL systems, which as a distributed modeling toolkits consists of server part and client part. In a general FL setting, the server part is the central aggregator installed in a trusted cloud environment, while the client part of software can operate in different operating environments on client devices. The server and clients are connected via Internet and typically with the help of Remote Procedure Call (RPC) interface for coordinating (Zhang et al., 2022f; Abadi et al., 2016; Liu et al., 2021b; Beutel et al., 2020; He et al., 2020b; Foley et al., 2022). We use four colors to represent the four FL roles and the colors with grid lines indicate non-essential roles. For example, in Fig. 2, the UCDA server takes on the roles of model user, coordinator and audior in traditional FL. However, it is no necessary to hold training data or validation data, so the role of data owner is non-essential. To illustrate the workflow of traditional FL, we leverage the vanilla FL framework Federated Averaging (FedAvg) (McMahan et al., 2017; Bonawitz et al., 2019) as an example.
@@
First, the FL server pre-defines the objective modeling task and initializes the server process. Secondly, the coordinator in server-side specifies a preset global model and the operational parameters. Thirdly, the coordinator discovers the availability of clients’ FL services, boardcasts the global model and training config to them. The training config contains bath size, local epoch round, optimizer parameters and so on. Then, the coordinator will wait for the trained results contributed by the coordinator in clients-side and drop those clients with network problems. Finally, the server aggregates the trained resultes received from various clients into the global model and begins a new round based on this aggregated global model. The aggregation strategy adopted in FedAvg is the weighted model parameters based on the size of local dataset, which means the global objective of FL can be regarded as a joint objective function of clients. By this way, the FL server can learn a generalized global model by jointly optimizing all local optimization objectives and incorporating the latent knowledge from the local models. Although the auditor component was not included in earlier FedAvg, it play an important role in the later business-ready FL frameworks (Liu et al., 2021b; Roth et al., 2022; Ziller et al., 2021).
@@
However, in comparing FedAvg workflow described above with Fig. 2, it is easy to notice that the client part has been excluded. This is because we are elaborating from a server-side perspective, which is usual way FL is presented (McMahan et al., 2017; Li et al., 2020c; Caldas et al., 2018). Actually, the underlying reason is that in traditional FL, the client-side process is tightly coupled with server-side process, and there is no alternative for clients other than to either accept or reject the training scheduling from the server wholesale. So the clients are not considered as an autonomous entities but rather work as subordinates to server. In this server-domianted cooperation framework, the benefits and autonomy of clients are compromised, which hinders their enthusiasm to participate in FL network and subsequently limits the applicability of FL. From this perspective, we summarize the limitations of traditional FL in the next section, which motivates us to explore more innovative sustainable FL cooperation frameworks.
3.2. Limitations of Traditional FL
@@
Previous surveys (Kairouz et al., 2021; Zhang et al., 2022b; Alazab et al., 2021; Nguyen et al., 2021; Zhu et al., 2022; Li et al., 2020e; Yang et al., 2019; Tan et al., 2022b) has extensively discussed the challenges in FL systems from various aspects. However, the cooperation mechanism of FL systems has been overlook because almost all mainsteam FL frameworks follow a same FL prototype (McMahan et al., 2017), which shape the form of current FL frameworks: a modeling software. We summarize three inherent limitations of traditional FL cooperation mechanism: (1) Server-client Coupling, (2) Low Model Reusability, (3) Non-public.
3.2.1. Server-client Coupling
The tightly-coupled server-client design is a major limitation of FL systems. From the perspective of FL service providers, adapting the programs to heterogeneous client hardware and software components, such as various operating and database systems, processor and storage architectures, communication protocols, energy constrains and data licenses, is a challenging task that significantly increases the complexity of the FL systems.
@@ 3.2.1. Server-client Coupling
On the other hand, the invasive software deploy mode compromises the integrity of client environments and expose them to new privacy risks. Specifically, the coordinator components (client-side) pushed by the server may not offer demanded privacy control mechanisms (Zeng et al., 2021; Caldas et al., 2018; McMahan et al., 2017), or cause resource depletion on client-side (Bonawitz et al., 2019; Niu et al., 2020; Chen et al., 2020), or even piggyback malicious executable codes (Li et al., 2017b). So the auditor role of client is non-essential as depicted in Fig. 2, not only because the client maybe lacks a corresponding policy for FL training, but also because its privacy is not completely under its control. Likewise, the malicious clients can also exploit the vulnerability in the aggreagation strategy to currupt the FL tranining process (Bouacida and Mohapatra, 2021; Sattler et al., 2020; Park et al., 2021; Fang et al., 2020) or insert backdoors (Bagdasaryan et al., 2020; Wang et al., 2020a). In addition, the unstable network environment can drive clients to drop out from traning (i.e. straggler effect), thereby reducing system efficiency (Reisizadeh et al., 2019; Park et al., 2021). Therefore, the server-client coupling design of traditional FL systems make them susceptible to unpredictable runtime environments, leading to system vulnerability and low reliability.
3.2.2. Low Model Reusability
@@
The traditional FL scheduling follows a task-centric manner and terminates once the training reaches a preset number of rounds or meets traget metrics on global model set by FL server (Bonawitz et al., 2019). As a result, only FL server can guarantee having the latest global model after the task is terminated. This ad-hoc modeling paradigm results in low model reusability and transportability. For example, if a client who participated in the previous training turn wants to continue training, they can only start the task from scratch unless they have the up-to-date global model. Meanwhile, since only FL server is able to maintain the complete modeling trajectory, it is difficult for the client to roll back the training itself to eliminate the potential privacy risk. Furthermore, the non-deliverable scheduling mechanism of FL tasks also hinders inter-task model reuse, which leads to unnecessary wasted energy and time on participants that have been involved in similar tasks.
3.2.3. Non-public
@@
As we mention in Section 2.1, except PySyft (Ziller et al., 2021), the application scenarios of mainstream FL frameworks (Liu et al., 2021b; Abadi et al., 2016; Zeng et al., 2021; Caldas et al., 2018; Ludwig et al., 2020; He et al., 2020b; Beutel et al., 2020; Roth et al., 2022) aim to provide private collaborative ML traning service, and there is no any accessible FL platform for the public. Although there have been real-world deployment practices of FL for the public with scales of millions (Bonawitz et al., 2019) and billions (Niu et al., 2020), these have been carried out only by tech giants with a massive base of active users. For an individual user, there is no practical way to organize such a large-scale FL training network.
@@
But in fact, due to the limitations in the cooperation mechanism mentioned above, data owners are not sufficiently motivated to participate in this server-take-all FL training network even if it is public accessible. Therefore, the cornerstone of buliding a sustainable open FL platform is to create a reciprocal FL cooperation framework, followed by corresponding multi-source knowledge aggregation strategies, which we discuss in the following sections.
4. Query-based Federated Learning
4.1. Overview
@@
Let us continue by establishing a sustainable open FL platform based on a query-based cooperation framework. An overview of this platform is presented in Fig. 3, the desin philosophy behind this framework is to break the coupling between FL server and clients. In the query-based FL systems, all traditional FL roles and components are maintained on an open model repository called Model Community. The Model Community privdes a one-stop ML models redistribution and reuse service, including model indexing, automatic batch model reuse, license management, privacy control and so on. In addition to large-scale pretrained models like BERT (Devlin et al., 2019), BLOOM (Scao et al., 2022) with great generalization abilities, we also encourage individuals to upload their task-specific models trained on limited domain data to boost the knowledge mining within models, aka model mining (You et al., 2021). The derivatives of model mining can learn representations from multiple domains, resulting in more promising performance that can be evaluated by platform users. Furthermore, the contributors can release models under applicable licenses, granting them distribution control and legal protection of their intellectual property (IP). In summary, the properties of query-based FL are: (1) Model Agnostic, as there are no restrictions on the types and architectures of the models uploaded by users; (2) Contactless, as communication channels need not be maintained; (3) Community-powered, whereby sharing models enrichs the entire community.
@@
Figure 3. An overview of query-based FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
[…]
@@
Actually, we aim to advocate a novel SaaS (Brereton et al., 1999) ML platform with automatic batch model reuse integrated, which has potential to leverage the transportability of models to address previously unexplored ML problems. Due to the high computational demands of deep learning, current ML platforms primarily concentrate on computing, for example, MaaS, MLaaS, FLaaS provide ML models deployment and development services to handle user-specified tasks (ref. Section 2.2). On the other hand, there are several ML platforms provide open model search and download services. So, can we leverage leverage off-the-shelf open model platforms to build a query-based FL system? Unfortunately, these platforms are designed solely for sharing and are no suitable for more advanced functionalities such as model ensemble (Jacobs et al., 1991) and knowledge distillation (Hinton et al., 2014), we will explain the reasons in the following section.
4.2. How to Query for Models
@@
To establish a query-based FL platform, the first thing that comes to mind is how to query for models. Unlike traditional ML model sharing repositories that mainly query for a specific model by name, it requires an efficiency approach to export a batch of target models that ready for ensemble or distillation. We summaried the filter conditions of existing DNNs sharing repositories in Table. 2. The prevailing method for querying models involves searching for the desired model by its name, datasets used, and the associated tasks. To illustrate, one might search for the model name GPT (Radford et al., 2019), models trained on the MNIST dataset (LeCun et al., 2010), or models capable of performing image segmentation tasks. However, this model retrieval method requires the users have a strong priori knowledge in data science, thus raising the barrier for knowledge mining within models. For example, there is no effective way to acquire a batch of image classfication models that contains the knowledge of lesser panda for further distillation. A compromise solution is to manually search the schema of each dataset one-by-one and subsequently search for models trained on those datasets.
@@
As shown in Table. 2, most DNNs repositories are simply list the description of input/output (e.g., NVIDIA NGC, OpenVINO) or even just present the source codes (e.g., Tensorflow Hub, Pytorch Hub). This lack of unified convention for model input/output poses a challenge for query-based FL. Additionally, most of DNNs repositories do not enable querying models by licenses, resulting in the cumbersome task of individually handling model licenses and ensuring compatibility among different licenses. Hence, it is imperative to reconsider the design of DNNs repositories to enable quick identification of readily reusable models for knowledge aggregation. We further suggest following filter conditions for query-based FL.
@@
Table 2. Filter conditions and characteristics of DNNs repositories. ✓: Supported, ✗ : Unsupported, ! : Information provided but unsearchable, listed in descending order by number of released models. (Accessed on January 17, 2024)
@@
Hugging Face [^5] https://huggingface.co ! 470,263
Model Zoo [^6] https://modelzoo.co/ 3,245
Tensorflow Hub [^7] https://tfhub.dev/ ! ! 2,186
NVIDIA NGC [^8] https://catalog.ngc.nvidia.com/models ! ! ! 680
OpenVINO [^9] https://docs.openvino.ai/latest/model_zoo.html ! ! ! 277
Pytorch Hub [^10] https://pytorch.org/hub/ ! ! 52
4.2.1. Data Description
Similar with the data heterogeous challenges in FL (Li et al., 2022c). The local datasets of contributors have varing quality and contain intractable biases, imbalances and noisies that can be attributed to the natural characteristics of demographic or improper data collection mechanisms (Dayan et al., 2021). Besides, label errors pervasive even in open datasets (Northcutt et al., 2021). So, in addition to searching for domain-specific datasets based on their data descriptions, we are also seeking such descriptions for the purpose of future traceability and debugging. The data description can consist of statistical analysis results or the visualization diagrams that used to profile the data distribution (Li et al., 2020a) and complementary provenance information.
4.2.2. Workflow and History
The process of building an ML model is iterative, involving repeated hyperparameter tuning and architecture exploration, resulting in abundant workflow and historical trajectory data. This information includes pipelines, model structures, hyperparameter values for pre-training and fine-tuning, test metrics, and results. These data can be useful in filtering models that meet specific requirements, such as those with data standardization in preprocessing or evaluated using mean average precision (mAP). Instead of manually saving and uploading the logs and configuration files, a more efficient method is to leverage ML workflow management tools (Vartak et al., 2016), such as MLflow  [^11] https://mlflow.org and Neptune  [^12] https://neptune.ai, to automatically track and store the ML workflow during model building process. This information can also assist in identifying potential model plagiarism within the model community. In addition, to ensure that the computational consumption of models is within budget, Deep Learning Profiler  [^13] https://docs.nvidia.com/deeplearning/frameworks/dlprof-user-guide/index.html can be leveraged to generate a report that displays the FLOPS and bandwidth requirements.
4.2.3. Software Dependency
ML models are software that depend on underlying ML libraries, so it is important to declare the dependencies of the model to analyze software compatibility between batches of models. For instance, resource-constrained devices may need to trim down the list of software-dependent libraries to meet limited storage space requirements (David et al., 2021). In some cases, contributed models may rely on other models as dependencies. For example, Fast R-CNN (Girshick, 2015) uses VGG16 (Simonyan and Zisserman, 2014) as its backbone. It is crucial to release this information for further model license compatibility analysis.
@@ 4.2.3. Software Dependency
The aforementioned filter conditions provide comprehensive coverage of the ML modeling process. However, there are additional requirements depending on the reuse mechanisms of the model retrieval side. For example, FedAvg (McMahan et al., 2017) aggregates the local models weights element-wise, which requires full access to the models. In contrast, MoE with a gating network (Jacobs et al., 1991) only ensembles a batch of model outputs, so the individual models can remain blackboxes in this scenario. So, in the context of software licenses or model licenses, the batch models reused by FedAvg should be released as source code, while those reused by MoE can be released as binary executable modules (e.g. dynamic linking). The above distinction is crucial for ensuring that model reuse results meet the legal framework, and this has been overlooked in traditional FL. We will expand on this topic in the following section.
4.3. How to Reuse Batch of Models

### 29 Feb 2024 v3

@@ III-A Definition
Federated Learning [19, 8] is a collaborative machine learning modeling paradigm that enables sharing and aggregation of knowledge from multiple sources while maintaining the confidentiality of source data. Generally, FL systems consist of two main entities in terms of task organization: the server and the participants.
@@
Furthermore, FL entities can also serve multiple functional roles to support advanced features such as privacy enhancement [17, 16, 36], participant scheduling [15, 37], model verification [38, 39] and incentive mechanisms [40]. Recall that there are four roles defined in the FL standard [19]:
@@
  • Model User. The FL model users can request for FL services and preset the targeted task, and then establish cooperation with participants who provide data. This role can leverage the benefits of collaborative training to improve the preformance of its objective models.
@@
  • Coordinator. The FL coordinators are responsible for providing FL services to all FL entities. This role involves setting up communication channels with entities, initializing the execution environment of participants [41], scheduling the training and aggregation workflows to improve system efficiency. Additionaly, the FL coordinator provides privacy control mechanisms [16, 30, 42] for model users and authorization verification for participants to maintain the security of FL systems. Furthermore, the coordinator can hold a validation dataset for evaluate the models contributed by participants or detect potential disturbances from Byzantine attacks [43].
@@
  • Data Owner. The FL data owners are knowledge contributors of FL systems, they collect and desentize raw data to maintain a local dataset for federated training. Although they have full authority over data processing and modeling, they cannot share the raw data due to privacy concerns. To address these concerns, de-identification [44] and differential privacy [45] techniques can be applied to meet privacy budgets as required by privacy policies.
@@
  • Auditor. The FL auditors are responsible for formulating privacy control policies and establishing supervisory mechanisms that ensure the training process is compliant with data protection regulations (e.g., HIPAA [44], GDPR [3]) and preventing potential privacy breaches for both model users and data owners. Especially in FL, the latent knowledge in models can potentially reveal the sensitive information of training data [46, 47, 48], making it crucial for auditors to scrutinize the model transmission [49, 50] and verify the ownership [38, 39].
@@
Figure 3: An overview of traditional FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
@@
Fig. 3 illustrates the typical architecture of traditional FL systems, which consists of server part and client part as a distributed modeling toolkits. In a general setting, the server part is the central aggregator installed in a trusted cloud environment, while the client part of software can operate in different operating environments on client devices. The server and clients are connected via Internet and typically with the help of Remote Procedure Call (RPC) interface for coordinating [51, 52, 53, 54, 55, 56]. We use four colors to represent the four FL roles and the colors with grid lines indicate non-essential roles. For example, in Fig. 3, the UCDA server takes on the roles of model user, coordinator and audior. However, there is no necessary to hold training or validation data, so the role of data owner is non-essential.
@@
To illustrate traditional FL workflow, we leverage the vanilla FL framework Federated Averaging (FedAvg) [8, 6] as an example. First, the FL server pre-defines the objective modeling task and initializes the server process. Secondly, the coordinator in server-side specifies a preset global model and the operational parameters. Thirdly, the coordinator discovers the availability of clients’ FL services, boardcasts the global model and training config to them. The training configuration contains bath size, local epoch round, optimizer parameters and so on. Then, the coordinator will wait for the trained results contributed by the coordinator in clients-side and drop those clients with network problems. Finally, the server aggregates the trained resultes received from various clients into the global model and begins a new round based on this aggregated global model. The aggregation strategy adopted in FedAvg is the weighted average of model parameters based on the size of the local dataset, which means the global objective of FL can be regarded as a joint objective function of clients. Although the auditor component was not included in prototype FedAvg, it plays an important role in the business-ready FL frameworks [53, 57, 58].
@@
However, in comparing FedAvg workflow described above with Fig. 3, it is easy to notice that the client part has been excluded. In traditional FL, the client-side process is tightly coupled with server-side process, leaving clients to either fully accept or reject the training schedule imposed by the server with compromised audit control. So the clients are not considered as an autonomous entities but rather work as subordinates to server. From this perspective, we summarize the limitations of traditional FL in the next section, which motivates us to explore more innovative sustainable FL cooperation frameworks.
III-B Limitations of Traditional FL
@@
Previous surveys [23, 59, 60, 33, 35, 4, 20, 27] has extensively discussed the challenges in FL systems from various aspects. However, the cooperation mechanism of FL systems has been overlook because almost all mainsteam FL frameworks follow a same prototype [8], which shapes the current FL form: a modeling software. We summarize three inherent limitations of traditional FL cooperation mechanisms: (1) Server-client Coupling, (2) Low Model Reusability, (3) Non-public.
III-B1 Server-client Coupling
@@
The invasive software deploy mode compromises the integrity of client environments and exposes them to new privacy risks. Specifically, the coordinator components (client-side) pushed by the server may not offer demanded privacy control mechanisms [61, 62, 8], or cause resource depletion on client-side [6, 36, 63], or even piggyback malicious executable codes [64]. So the auditor role of client is non-essential as depicted in Fig. 3, not only because the client maybe lacks a corresponding policy for FL training, but also because its privacy is not completely under its control. Likewise, the malicious clients can also exploit the vulnerability in the aggreagation strategy to currupt the FL tranining process [65, 43, 66, 67] or insert backdoors [68, 69]. In addition, the unstable network environment can drive clients to drop out from traning (i.e., straggler effect), thereby reducing system efficiency [70, 66]. Therefore, the server-client coupling design of traditional FL systems make them susceptible to unpredictable runtime environments, leading to system vulnerability and low reliability.
III-B2 Low Model Reusability
@@
The traditional FL scheduling follows a task-centric manner and terminates once the training reaches a preset number of rounds or meets traget metrics on global model set by FL server [6]. As a result, only FL server can guarantee having the latest global model after the task is terminated. This ad-hoc modeling paradigm results in low model reusability and transportability. Since only FL server is able to maintain the complete modeling trajectory, it is difficult for the client to roll back the training itself to eliminate the potential privacy risk. Meanwhile, the non-deliverable tasks scheduling mechanism also hinders inter-task model reuse.
III-B3 Non-public
@@
Except PySyft [58], the application scenarios of mainstream FL frameworks [53, 52, 61, 62, 71, 55, 54, 57] aim to provide private collaborative ML traning service, and there is no any accessible FL platform for the public‡I-A. Although there have been real-world deployment practices of FL for the public with scales of millions [6] and billions [36], these have been carried out only by tech giants with a massive base of active users. For an individual user, there is no practical way to organize such a large-scale FL training network.
@@
Due to the limitations in the cooperation mechanism mentioned above, data owners are not sufficiently motivated to participate in this server-take-all FL training network even if it is public accessible. Therefore, the cornerstone of buliding a sustainable open FL platform is to create a reciprocal FL cooperation framework, followed by corresponding multi-source knowledge aggregation strategies, which we survey through the following two themes:
In the query-based FL platform, which is an open model sharing and reusing platform empowered by the community for model mining, we explore a wide range of valuable topics, including the availability of up-to-date model repositories for model querying, legal compliance analysis between different model licenses, and copyright issues and IP protection in model reusing. In the contract-based FL platform, which is an opt-in model training network that allows multi-round communication and monetization, we explain how to select training frameworks according to microtasks and discuss the pros and cons of Web3-based methods. We will detail discuss these topics in the following sections.
IV Query-based Federated Learning
IV-A Overview
@@
Let us continue by establishing a sustainable open FL platform based on a query-based cooperation framework. An overview of this platform is presented in Fig. 4, the desin philosophy behind this framework is to break the coupling between FL server and clients. In the query-based FL systems, all traditional FL roles and components are maintained on an open model repository called Model Community. The Model Community privdes a one-stop ML models redistribution and reuse service, including model indexing, automatic batch model reuse, license management, privacy control and so on. In addition to large-scale pretrained models like BERT [72], BLOOM [73] with great generalization abilities, we also encourage individuals to upload their task-specific models trained on limited domain data to boost the knowledge mining within models, aka model mining [74]. The derivatives of model mining can learn representations from multiple domains, resulting in more promising performance that can be evaluated by platform users. Furthermore, the contributors can release models under applicable licenses, granting them distribution control and legal protection of their intellectual property (IP). In summary, the properties of query-based FL are: (1) Model Agnostic, as there are no restrictions on the types and architectures of the models uploaded by users; (2) Contactless, as communication channels need not be maintained; (3) Community-powered, whereby sharing models enrichs the entire community.
@@
Figure 4: An overview of query-based FL systems. (U: model User, C: Coordinator, D: Data owner, A: Auditor)
@@
Actually, we aim to advocate a novel Software-as-a-Service (SaaS) [75] ML platform with automatic batch model reuse integrated, which has potential to leverage the transportability of models to address previously unexplored ML problems. Due to the high computational demands of deep learning, current ML platforms primarily concentrate on computing, for example, MaaS, MLaaS, FLaaS provide ML models deployment and development services to handle user-specified tasks‡I-B. On the other hand, there are several ML platforms provide open model search and download services. So, can we leverage leverage off-the-shelf public model repositories to build a query-based FL system? Unfortunately, these sites are designed solely for sharing and are not suitable for more advanced functionalities such as model ensemble [76] and knowledge distillation (KD) [77], and we will explain the reasons in the following section.
IV-B How to Query for Models
@@
To establish a query-based FL platform, the first thing that comes to mind is how to query for models. Unlike traditional ML model sharing repositories that mainly query for a specific model by name, it requires an efficiency approach to export a batch of target models that are ready for ensemble or distillation. We summarized the filter conditions of existing DNNs sharing repositories [^1] Hugging Face: https://huggingface.co; Model Zoo: https://modelzoo.co; Tensorflow Hub: https://tfhub.dev; NVIDIA NGC: https://catalog.ngc.nvidia.com/models; OpenVINO: https://docs.openvino.ai/latest/model_zoo.html; Pytorch Hub: https://pytorch.org/hub in TABLE I. The prevailing method for querying models involves searching for the desired model by its name, used datasets, and the associated tasks. For instance, one might search for the model name GPT [78], models trained on the MNIST dataset [79], or models capable of performing image segmentation tasks. However, this model retrieval method requires the users have a strong priori knowledge in data science, thus raising the barrier for knowledge mining within models. As an example, there is no effective way to acquire a batch of image classfication models that contains the knowledge of lesser panda for further distillation. A compromise solution is to manually search the schema of each dataset one-by-one and subsequently search for models trained on those datasets.
@@
As shown in TABLE I, most DNNs repositories are simply listing the description of input/output (e.g., NVIDIA NGC, OpenVINO) or even just present the source codes (e.g., Tensorflow Hub, Pytorch Hub). This lack of unified convention for model input/output poses a challenge for query-based FL. Additionally, most of DNNs repositories do not enable querying models by licenses, resulting in the cumbersome task of individually handling model licenses and ensuring compatibility among different licenses. Hence, it is imperative to reconsider the design of DNNs repositories to enable quick identification of readily reusable models for knowledge aggregation. We further suggest following filter conditions for query-based FL: 1) Data Description; 2) Workflow and History; 3) Software Dependency; 4) Fairness and Robustness‡II.
@@
TABLE I: Filter conditions and characteristics of DNNs repositories. ✓: Supported, ✗: Unsupported, ! : Information provided but unsearchable, listed in descending order by number of released models. (Accessed on January 17, 2024)
@@
Hugging Face ! 470,263
Model Zoo 3,245
Tensorflow Hub ! ! 2,186
NVIDIA NGC ! ! ! 680
OpenVINO ! ! ! 277
Pytorch Hub ! ! 52
@@
The aforementioned filter conditions provide comprehensive coverage of the ML modeling process. However, there are additional requirements depending on the reuse mechanisms in the model retrieval side. For example, FedAvg [8] aggregates the local models weights element-wise, which requires full access to the models. In contrast, MoE [76] only ensembles a batch of model outputs, so the individual models can remain blackboxes in this scenario. So, in the context of software licenses or model licenses, the batch models reused by FedAvg should be released as source code, while those reused by MoE can be released as binary executable modules (e.g., dynamic linking). The above distinction is crucial for ensuring that model reuse results meet the legal framework, and this has been overlooked in traditional FL. We will expand on this topic in the following section.
IV-C Legal Considerations in Batch Model Reuse